OpenStudy (anonymous):

In a 45°- 45°- 90° right triangle, the length of the hypotenuse is 15. How long are the legs?

OpenStudy (anonymous):

@tcarroll010

OpenStudy (anonymous):

Because the legs will be equal in a 45-45-90 right triangle, the equation: a^2 + b^2 = c^2 becomes a^2 + a^2 = c^2 or, better, yet: 2a^2 = c^2 c=15, so: 2a^2 = 15^2 so, a = sqrt[(15^2)/2]

OpenStudy (anonymous):

so whats the othervall answer ?

OpenStudy (anonymous):

You first square 15. Take that # and divide by 2. Take that answer and get the square root of it.

OpenStudy (aroub):

Or, you can say that: in a 45-45-90 triangle hypotenuse=sqrt2 times any of the legs so now you just substitute.